# When Text Embedding Meets Large Language Model: A Comprehensive Survey

### v1

@@ III-B Data Annotation with LLMs
III-B1 Training Data Supervision
Compared to directly generating text from LLMs to construct training data, mining training data from large corpus is more effective in meeting the demand for diversity. As introduced in § III-A3, hard negatives are critical for text embedding models. Therefore, some studies leverage LLMs for hard negative mining: NV-Retriever [99] uses LLMs as embedding model and employs proposed “positive-aware mining methods” to identify hard negatives; Gecko [101] first retrieves relevant texts using a embedding model, then estimates the similarity between texts using LLMs to mine hard negatives. In addition to traditional training data composed of positive and negative samples, the application of LLMs enables the construction of training data with finer-grained similarity patterns: CLAIF and CLHALF [91] use LLMs to assign similarity scores to generated and existing text pairs, then train models using Mean Squared Error (MSE) loss and a proposed soft InfoNCE loss; NGCSE [92] employs LLMs to construct text pairs with similarity levels ranging from high to low, which equals to labeling text pairs with similarity scores. These pairs are then used for training with a proposed Hierarchical Triplet loss.
III-B2 Training Data Filtering
@@ III-B3 Clustering Supervision
TABLE IV: The Overview of LLMs as Text Encoders. The paradigm part is shortened as follows: supervised contrastive learning (SCL), weak supervised contrastive learning (WCL), unsupervised contrastive learning (UCL), instruction tuning (IT), prompt engineering (PE), next-token prediction (NTP), Generation-augmented Representation Learning (GRL), FLOPS regularization (FLOPs) and masked next-token prediction (MNTP).
@@
LLM (Encoder) Attention Projector Pooling PEFT Paradigm Instruction Setting Task
Sentence-T5 [113] T5 Frist / Mean Bi $\surd$ $\times$ WCL→SCL $\times$ ZS / FT STS / Clf.
PromptEOL [114] OPT, LLaMA Last Casual $\times$ $\times$ PE $\times$ ZS / ICL STS / Clf.
PromptEOL+CSE [114] OPT, LLaMA Last Casual $\times$ LoRA PE+SCL $\times$ ZS
MetaEOL [115] LLaMA, Mistral Last Casual $\times$ $\times$ PE $\times$ ZS / FT STS / Clf.
PromptSTH/SUM [116] OPT, LLaMA, Mistral Last Casual $\times$ $\times$ PE $\times$ ZS STS
BeLLM [117] LLaMA Last Bi $\times$ LoRA PE+SCL $\times$ ZS / FT STS / Clf.
GTR [118] T5 Mean Bi $\surd$ $\times$ WCL→SCL $\times$ ZS / FT IR
SGPT-IR [119] GPT-Neo Weighted Mean Casual $\times$ BitFit SCL $\times$ ZS / FT IR
@@
Llama2Vec [120] LLaMA Special Last Casual $\times$ LoRA SCL $\times$ FT IR
RepLLaMA [121] LLaMA Special Last Casual $\times$ LoRA SCL $\times$ ZS / FT IR
@@
BMRetriever [122] Pythia, Gemma, BioMistral Special Last Casual $\times$ LoRA WCL→SCL $\surd$ ZS / FT IR
@@
ChatRetriever [123] Qwen Special Seq Last Customized $\times$ LoRA SCL+MLM $\surd$ ZS / FT IR
@@
PromptReps [124] Mistral, Phi-3, LLaMA Last Casual $\surd$ $\times$ PE $\surd$ ZS IR
LMORT [125] GPT2, GPT-J Mean Casual $\surd$ $\times$ SCL $\times$ FT IR
@@
Mistral-SPLADE [126] Mistral Partial Mean Casual $\surd$ QLoRA PE+SCL+FLOPS $\times$ ZS IR
NV-Retriever [99] Mistral Mean Bi $\times$ LoRA SCL→SCL $\surd$ ZS IR
FollowIR [103] Mistral - Casual $\times$ LoRA - $\surd$ FT IR
@@
Promptriever [100] LLaMA Special Last Casual $\times$ LoRA SCL $\surd$ ZS / FT IR
@@
RARe [127] LLaMA, Mistral Mean / Last Bi / Casual $\times$ LoRA ICL+SCL $\surd$ ICL IR
@@
Echo [128] Mistral Partial Mean Casual $\times$ LoRA PE $\surd$ ZS Uni.
GenEoL [129] LLaMA, Mistral Mean Casual $\times$ $\times$ PE $\times$ ZS Uni.
E5-Mistral [102] Mistral Last Casual $\times$ LoRA SCL $\surd$ ZS Uni.
LLM2Vec [130] LLaMA, Mistral Mean Bi $\times$ LoRA MNTP→UCL $\surd$ ZS Uni.
Cpt [131] Unknown Special Last Unknown $\times$ $\times$ SCL $\times$ ZS Uni.
UDEVER [132] BLOOM Special Last Casual $\times$ BitFit SCL $\times$ ZS Uni.
InstructOR [104] GTR Mean Bi $\surd$ $\times$ SCL $\surd$ ZS Uni.
InBedder [133] OPT, LLaMA Last Casual $\times$ $\times$ PE+NTP $\surd$ ZS Uni.
GritLM [134] Mistral Mean Customized $\surd$ / $\times$ $\times$ SCL+NTP $\surd$ ZS / FS Uni.
GTE-Qwen1.5 [135] Qwen Last Bi $\times$ $\times$ WCL→SCL $\surd$ ZS Uni.
GTE-Qwen2 [135] Qwen Last Bi $\times$ $\times$ WCL→SCL $\surd$ ZS Uni.
NV-Embed [136] Mistral Mean Bi $\surd$ LoRA WCL→SCL $\surd$ ZS Uni.
ULLME [137] LLaMA, Mistral, Phi Mean Bi / Casual $\times$ LoRA GRL $\times$ ZS Uni.
L3Prune [138] LLaMA, Mistral, Qwen, Phi Weighted Mean Casual $\times$ LoRA SCL $\surd$ ZS Uni.
BGE-ICL Mistral, Gemma Special Last Bi / Casual $\times$ LoRA ICL+SCL $\surd$ ZS / ICL Uni.
@@
MoEE [139] Deepseek, Qwen, OLMoE Last Casual $\times$ $\times$ PE $\times$ ZS Uni.
IV LLMs as Text Embedder
LLMs commonly adopt encoder-decoder architecture, e.g., T5 [140], or decoder-only architecture, e.g., GPT-3 [141]. After going through pre-training, instruction tuning, and alignment with human preferences [142], LLMs can perform a wide range of tasks. However, these steps do not support obtaining high-quality embeddings from LLMs. To address this issue, exploring methods to obtain representations from LLMs directly is a promising direction. In Table IV, we list 35 LLM-based Embedder and their detailed information. Specifically, we focus on their backbone selection, architectural improvement, training methods, and evaluation protocols.
IV-A Backbone Selection
The LLM-based embedders appearing in Table IV use numerous open-source LLMs as backbones. Among Encoder-Decoder backbones, T5 (GTR) has a dominant position; Among decoder-only backbones, the most popular is Mistral (18 / 32), followed by LLaMAc (15 / 32) and OPT (4 / 32). The dominance of T5 is not surprising, as T5 and its variants have been leading innovations in encoder-decoder PLMs. Meanwhile, the popularity of the Mistral is strongly connected to its excellent performance. LLM2Vec [130] find that Mistral’s embedding performance will not drop significantly compared to the other LLMs when converting casual attention to bi-directional attention. We speculate that this characteristic should derive from its pre-training or instruction fine-tuning phase, but no information has been revealed.
IV-B Architecture Improvement
Denoted $d$ is the dimension of hidden states, $L$ is the number of transformer layers, and $c$ is the vocabulary size. Without loss of generality, we divide the LLM $F$ into multiple Transformer layers and a decoding layer, which can be expressed as
$F=g\circ[f^{(L)}\circ\cdots\circ f^{(1)}]\circ f^{(0)}$ (17)
IV-B1 Pooling Strategy
As with traditional pre-trained language models, there is a decoding layer within the LLM to complete the mapping from the hidden state space to the token space. To the best of our knowledge, almost all LLMs follow the design of GPT [143], using a simple unbiased linear layer as the decoding layer. When LLMs are used as textual embedders, embeddings are obtained by a specific pooling strategy at hidden states, while mapping to token space is no longer necessary. Therefore, most methods directly discard the decoding layer and opt for a pooling strategy on the hidden state of the last transformer layer’s output to obtain a single embedding for each text.
$[{\bm{h}}^{(L)}_{1},\cdots{\bm{h}}^{(L)}_{n}]=\mathcal{H}=f(x),\quad x\in\mathcal{X},$ (18)
[…]
where $n$ is the token number of the input $x$ depending on the tokenizer’s vocabulary of the LLMs, and $\{\alpha_{i}\}_{i=1}^{n}$ is the weight satisfying $\sum_{i}^{n}\alpha_{i}=1$.
@@ IV-B1 Pooling Strategy
First pooling (only $\alpha_{1}=1$) is a common way of obtaining embedding for PLMs that use bi-directional attention. For example, BERT splices [CLS] before each text, and the embedding output from the [CLS]’s position will be used as the embedding for the whole text [105]. Although T5-like models do not have the special token like [CLS] in BERT, Sentence-T5 [144] use the first embedding output by T5’s encoder and the [START] embedding output by T5’s decoder as the embedding of the whole text; however, these pooling strategies obtain worse performance than mean pooling. For the decoder-only LLMs with casual attention, the first pooling obviously does not work, as the embedding from the first position can’t be included in the semantics of the subsequent content.
(Weighted) Mean Pooling
Mean pooling ($\alpha_{i}=\frac{1}{n}$ for each $i$) averages the Embedding of each position and shows better performance in BERT and T5 than first pooling. Meanwhile, in decoder-based LLMs, it is not reasonable to assign the same weight to each position because the latter position will be able to see more semantic information due to casual attention [119]. Therefore, SGPT [119] suggests the use of weighted mean pooling ($\alpha_{i}=\frac{i}{\sum_{i=1}^{n}i}$) with the intuition that the latter position should be given larger weight. L3Prune [138] follows SGPT’s pooling strategy to fine-tune the pruned LLM-based embedder.
(Special) Last Pooling
Last pooling (only $\alpha_{n}=1$) is a new pooling strategy emerging in the LLM era to ensure that embedding can contain information about the whole text. However, decoder-only LLMs are trained by next token prediction; the embedding at the last position will align with the potential next token embedding [114]. Without additional interventions, there is no guarantee that the embedding contains the semantics of the entire text. Two common interventions here are (1) using prompts to induce the model to summarize the semantics of the whole text in the last position and (2) introducing special tokens and incrementally fine-tuning them. For the first line of works [114, 116, 117, 124, 129, 139], a prompt template such as The sentence [X] means in a word:’’ is introduced, where [X] is a placeholder. In practice, [X] is filled with the input text, and the last pooling strategy is used to obtain the text embedding. PromptEOL [114] first proposed the idea, following up on the success of PromptBERT [145]. The templates generate text embeddings that perform better on the downstream task when it is related to the downstream task [115] or when CoT [116] is introduced. Still, the core idea in the line of work is consistent, i.e., convert text embedding to language modeling and guide the model to summarize the whole text semantics in the last position. These prompt-based methods, although able to stimulate the potential of decoder-only LLM without training, bring additional computational overhead because the same template content is always encoded over and over again. If we have some fine-tuning data, then the model can be guided to summarize the semantics at the location of the special token without the need for a prompt template. For the second line of works [131, 132, 120, 124, 123, 122, 121], a special token, such as <EOS>, is added to the end of the text, and the special token’s hidden states output by the encoder $f$ is regarded as the whole text’s embedding. These methods are marked as “special last” in the Table IV. A special case is ChatRetriever [123], which adds a token sequence $[\texttt{EMB}_{1}],\cdots[\texttt{EMB}_{t}]$ at the end of the input text. The authors argue that these $t$ consecutive special tokens serve as a chain of thought that extends and guides the learning space for more effective representations.
@@
Echo [128] propose a special prompt: Rewrite the sentence: [x], rewritten sentence: [x], where [x] is the placeholder. In practice, both placeholders are filled with the same text, and the mean pooling strategy is used to obtain the text embedding, but it is pooled only within the range of the second occurrence of the text. In this way, the entire text is already present in the first placeholder, so each token populated in the second placeholder can access the entire text information through casual attention. Mistral-SPLADE [126] follows Echo’s approach and uses the obtained embedding to generate sparse representations.
Trainable Pooling
@@
Some works add complex modules with attention mechanisms before conventional mean-pooling. NV-Embeder [136] introduces an additional latent attention layer before mean pooling to post-interact with the hidden states $\mathcal{H}$. The latent attention layer consists of a cross-attention block and an MLP, where the key and value matrix used in the cross-attention are learnable and low-rank, and $\mathcal{H}$ is considered as the query. Inspired by the success of alignment and uniformity in visual representation learning [146], [125] uses the probe dataset metrics to measure alignment and uniformity of each LLM layer’s output hidden layer states and collect the hidden states from the layer with optimal alignment and uniformity, which denotes $\mathcal{H}^{a}$ and $\mathcal{H}^{u}$ separately. Then, a multi-layer projector named LMORT is introduced to fuse the information in both $\mathcal{H}^{a}$ and $\mathcal{H}^{u}$, where each layer consists of a bi-directional self-attention block, a bi-directional cross-attention block, a feed-forward block, and inter-block residual mechanism. Experiments show that the performance of retrieval tasks can be improved by only fine-tuning the LMORT’s parameters. Similarly, [147] proposes a multi-layer trainable pooling method, which fully utilizes the hidden states from multiple layers. Specifically, the text embeddings are obtained from each layer with the same pooling strategy and combined into a matrix $\mathcal{H}^{L}\in\mathbb{R}^{l\times d}$. Then, a pooling layer is introduced, which consists of a cross-attention block, a feed-forward block, and the inter-block residual mechanism. When calculating cross-attention, $\mathcal{H}^{L}$ and a learnable weight matrix $\mathcal{W}$ are summed first and multiplied with $W_{k}$ and $W_{v}$, respectively, to get the key and value. At the same time, another learnable matrix $Q$ is treated as a fixed query among different input texts.
IV-B2 Attention Mechanism

### v2

@@ 3.2 Data Annotation with LLMs
3.2.1 Training Data Supervision
Compared to directly generating text from LLMs to construct training data, mining training data from large corpus is more effective in meeting the demand for diversity. As introduced in § 3.1.3, hard negatives are critical for text embedding models. Therefore, some studies leverage LLMs for hard negative mining: NV-Retriever [149] uses LLMs as embedding model and employs proposed “positive-aware mining methods” to identify hard negatives; Gecko [152] first retrieves relevant texts using a embedding model, then estimates the similarity between texts using LLMs to mine hard negatives. In addition to traditional training data composed of positive and negative samples, the application of LLMs enables the construction of training data with finer-grained similarity patterns: CLAIF and CLHALF [142] use LLMs to assign similarity scores to generated and existing text pairs, then train models using Mean Squared Error (MSE) loss and a proposed soft InfoNCE loss; NGCSE [143] employs LLMs to construct text pairs with similarity levels ranging from high to low, which equals to labeling text pairs with similarity scores. These pairs are then used for training with a proposed Hierarchical Triplet loss.
3.2.2 Training Data Filtering
@@ 3.2.3 Clustering Supervision
4 LLMs as Text Embedder
LLMs commonly adopt encoder-decoder architecture, e.g., T5 [94], or decoder-only architecture, e.g., GPT-3 [162]. After going through pre-training, instruction tuning, and alignment with human preferences [163], LLMs can perform a wide range of tasks. However, these steps do not support obtaining high-quality embeddings from LLMs. To address this issue, exploring methods to obtain text embeddings from LLMs directly is a promising direction. In Table 4, we list 37 LLM-based Embedders and their detailed information. Specifically, we focus on their backbone selection, architectural improvement, training methods, and evaluation protocols.
4.1 Backbone Selection
Model
The LLM-based embedders appearing in Table 4 use numerous open-source LLMs as backbones. Among the encoder-decoder backbones, T5 (GTR) has a dominant position, which is not surprising, as T5 and its variants have been leading innovations in encoder-decoder PLMs. Among decoder-only backbones, the most popular is Mistral (17 / 37), followed by LLaMA (15 / 37) and Qwen (5 / 37). The popularity of the Mistral is strongly connected to its excellent performance: a recent study [164] find that Mistral’s embedding performance will not drop significantly compared to the other LLMs when converting causal attention to bi-directional attention. This characteristic may derive from its pre-training or instruction fine-tuning phase, but no information has been revealed.
Size
Interestingly, the only several attempts to use LLM with larger LLMs lead to even worse performance [165, 166].
4.2 Architecture Improvement
Denoted $d$ is the dimension of hidden states, $L$ is the number of transformer layers, and $c$ is the vocabulary size. Without loss of generality, we divide the LLM $F$ into multiple Transformer layers and a decoding layer, which is denoted as
$F=g\circ[f^{(L)}\circ\cdots\circ f^{(1)}]\circ f^{(0)}=g\circ f$
4.2.1 Pooling Strategy
Similar to traditional language models, there is a decoding layer within the LLM to complete the mapping from the hidden state space to the token vocabulary space. To the best of our knowledge, almost all LLMs follow the original design of GPT [167], using a simple unbiased linear layer as the decoding layer. When LLMs are used as textual embedders, embeddings are obtained by a specific pooling strategy at hidden states, while mapping to token space is no longer necessary. Therefore, most methods directly discard the decoding layer and opt for a pooling strategy on the hidden state of the last transformer layer’s output to obtain a single embedding for each text.
$[{\bm{h}}^{(L)}_{1},\cdots{\bm{h}}^{(L)}_{n}]=\mathcal{H}=f(x),\quad x\in\mathcal{X},$
[…]
where $n$ is the token number of the input $x$ depended on the tokenizer of the LLMs, and $\{\alpha_{i}\}_{i=1}^{n}$ is the weight satisfying $\sum_{i}^{n}\alpha_{i}=1$.
Four pooling methods fit this scenario: (Weighted) Mean Pooling, First Pooling, Last Pooling and Partial Pooling. However, some methods are not simply weighting the hidden states of the last layer. For example, some methods use learnable modules to interact with the hidden states of the last layer (Post-Interaction Pooling), while others try to augment the final embedding with other information (e.g., attentional weights, hidden states, etc.) that utilizes the output of the middle layer (Multi-Layer Pooling). We cover them one by one below.
(Weighted) Mean Pooling
Mean pooling ($\alpha_{i}=\frac{1}{n}$ for each $i$) averages the embedding of each position and shows better performance in BERT and T5 than first pooling [72, 168]. Meanwhile, in decoder-based LLMs, it is not reasonable to assign the same weight to each position because the latter position will be able to see more semantic information due to causal attention [169]. Therefore, SGPT [169] suggests the use of weighted mean pooling ($\alpha_{i}=\frac{i}{\sum_{i=1}^{n}i}$) with the intuition that the latter position should be given larger weight. L3Prune [170] follows SGPT’s pooling strategy to fine-tune the pruned LLM-based embedders.
Table 4: The Overview of LLMs as Text Encoders. The paradigm part is shortened as follows: supervised contrastive learning (SCL), weak supervised contrastive learning (WCL), unsupervised contrastive learning (UCL), training-free (TF), next-token prediction (NTP), Generation-augmented Representation Learning (GRL), FLOPS regularization (FLOPs), masked next-token prediction (MNTP) and In-Context Learning (ICL).
@@ 4.2.1 Pooling Strategy
LLM (Encoder) Pooling Attention Projector PEFT Input Format Paradigm Setting Task
Sentence-T5 [168] T5 Frist / Mean Bi $\surd$ $\times$ - WCL→SCL ZS / FT STS / Clf.
PromptEOL [166] OPT, LLaMA Last Causal $\times$ $\times$ Prompt TF ZS / ICL STS / Clf.
MetaEOL [171] LLaMA, Mistral Last Causal $\times$ $\times$ Prompt TF ZS / FT STS / Clf.
PromptSTH/SUM [172] OPT, LLaMA, Mistral Last Causal $\times$ $\times$ Prompt TF ZS STS
Token Prepending [173] LLaMA, Qwen, Gemma Last Causal $\times$ $\times$ - TF ZS / FT STS / Clf.
BeLLM [174] LLaMA Last Bi $\times$ LoRA - SCL ZS / FT STS / Clf.
GTR [175] T5 Mean Bi $\surd$ $\times$ - WCL→SCL ZS / FT IR
SGPT-IR [169] GPT-Neo Weighted Mean Causal $\times$ BitFit - SCL ZS / FT IR
@@
Llama2Vec [176] LLaMA Special Last Causal $\times$ LoRA - EBAE+EBAR→SCL FT IR
RepLLaMA [177] LLaMA Special Last Causal $\times$ LoRA - SCL ZS / FT IR
@@
BMRetriever [178] Pythia, Gemma, BioMistral Special Last Causal $\times$ LoRA Instruction WCL→SCL ZS / FT IR
@@
ChatRetriever [179] Qwen Special Seq Last Customized $\times$ LoRA Instruction SCL+MLM ZS / FT IR
@@
PromptReps [180] Mistral, Phi-3, LLaMA Last Causal $\surd$ $\times$ Instruction+Prompt TF ZS IR
LMORT [181] GPT2, GPT-J Multi-Layer Causal $\times$ $\times$ - SCL FT IR
@@
Mistral-SPLADE [182] Mistral Post Causal $\times$ QLoRA Prompt SCL+FLOPS ZS IR
NV-Retriever [149] Mistral Mean Bi $\times$ LoRA Instruction SCL→SCL ZS IR
@@
Promptriever [151] LLaMA Special Last Causal $\times$ LoRA Instruction SCL ZS / FT IR
@@
RARe [183] LLaMA, Mistral Mean / Last Bi / Causal $\times$ LoRA Instruction+Examples SCL ICL IR
DEBATER [184] MiniCPM Special Seq Last Causal $\times$ LoRA Instruction SCL+SD ZS IR
@@
Echo [185] Mistral Partial Mean Causal $\times$ LoRA Instruction+Prompt TF ZS Uni.
GenEoL [186] Mistral Mean Causal $\times$ $\times$ Prompt TF ZS Uni.
@@
MoEE [187] Deepseek, Qwen, OLMoE Multi-Layer Causal $\times$ $\times$ Prompt TF ZS Uni.
ReBA [188] GPT-2, LLaMA Multi-Layer Causal $\times$ $\times$ - TF ZS Uni.
E5-Mistral [153] Mistral Last Causal $\times$ LoRA Instruction SCL ZS Uni.
LLM2Vec [164] LLaMA, Mistral Mean Bi $\times$ LoRA Instruction MNTP→UCL ZS Uni.
Cpt [189] Unknown Special Last Unknown $\times$ $\times$ - SCL ZS Uni.
UDEVER [190] BLOOM Special Last Causal $\times$ BitFit - SCL ZS Uni.
InstructOR [67] GTR Mean Bi $\surd$ $\times$ Instruction SCL ZS Uni.
InBedder [191] OPT, LLaMA Last Causal $\times$ $\times$ Instruction+Prompt NTP ZS Uni.
GritLM [165] Mistral Mean Customized $\surd$ / $\times$ $\times$ Instruction SCL+NTP ZS / FS Uni.
GTE-Qwen2 [192] Qwen Last Bi $\times$ $\times$ Instruction WCL→SCL ZS Uni.
NV-Embed [193] Mistral Post Bi $\times$ LoRA Instruction WCL→SCL ZS Uni.
MLTP [194] Mistral Multi-Layer Bi / Causal $\times$ $\times$ Instruction SCL ZS Uni.
ULLME [195] LLaMA, Mistral, Phi Mean Bi / Causal $\times$ LoRA - GRL ZS Uni.
L3Prune [170] LLaMA, Mistral, Qwen, Phi Weighted Mean Causal $\times$ LoRA Instruction SCL ZS Uni.
BGE-ICL [196] Mistral, Gemma Special Last Bi / Causal $\times$ LoRA Instruction+Examples SCL ZS / ICL Uni.
Linq-Embed [197] Mistral Last Causal $\times$ LoRA Instruction SCL ZS Uni.
@@
First pooling (only $\alpha_{1}=1$) is a common way of obtaining embedding for PLMs with bi-directional attention. For example, BERT splices the special token [CLS] before the input text, and the embedding output from the [CLS]’s position will be used as the embedding for the whole text [72]. Although T5-like models do not have the special token like [CLS] in BERT, Sentence-T5 [168] use the first embedding output by T5’s encoder and the [START] embedding output by T5’s decoder as the embedding of the whole text; however, these pooling strategies obtain slightly worse performance than mean pooling. For the decoder-only LLMs with causal attention, the first pooling does not work, as the embedding of the first position can’t be included in the semantics of the subsequent content.
Last Pooling
Last pooling (only $\alpha_{n}=1$) is a new pooling strategy emerging in the LLM era. Due to the uni-directionality of causal attention, only the last position shows information about the whole text. However, decoder-only LLMs are pre-trained by the next token prediction, and the embedding at the last position will align with the potential next token embedding [166]. Without additional interventions, there is no guarantee that the embedding contains the semantics of the entire text. Three common interventions here are
  • Prompt-Based Last Pooling. A special prompt is introduced to induce the model to summarize the semantics of the whole text in the last position [166, 172, 174, 180, 186, 187]. For example, PromptEOL [166] propose a prompt template The sentence [X] means in a word:‘‘, where [X] is a placeholder. In practice, [X] is filled with the input text, and the last pooling strategy is used to obtain the text embedding. Although there are many variants, the core idea in the line of work is consistent, i.e., convert text embedding to language modeling and guide the model to summarize the whole text semantics in the last position. These methods are marked as “last” in the “pooling” column of Table 4.
  • Special-Token Last Pooling Some works [189, 190, 176, 180, 179, 178, 177] introduce a special token, such as <EOS>, to obtain embedding in the last position, while the LLM is fine-tuned incrementally, learning to converge semantics at the position of the special token. These methods are marked as “special last” in the “pooling” column of Table 4.
  • Special-Sequence Last Pooling ChatRetriever [179] and DEBATER [184] add a special token sequence, such as $[\texttt{EMB}_{1}],\cdots,[\texttt{EMB}_{t}]$, at the end of the input text. The authors argue that these $t$ consecutive special tokens serve as a chain of thought that extends and guides the learning space for more effective embeddings. These methods are marked as “special seq last” in the “pooling” column of Table 4.
@@
Partial Pooling ($\alpha_{i}=1$ for $i$ is greater than a specific $k$) is also an effective method to mitigate the effects of causal attention, which requires prompt-based assistance. Echo [185] propose a prompt: Rewrite the sentence:[x], rewritten sentence:[x], where [x] is the placeholder. In practice, both placeholders are filled with the same text, and the mean pooling strategy is used to obtain the text embedding, but it is pooled only within the range of the second occurrence of the text. In this way, the entire text is already present in the first placeholder, so each token populated in the second placeholder can access the entire text information through causal attention. Mistral-SPLADE [182] follows Echo’s approach and uses the obtained embedding to generate sparse representations.
Post-Interaction Pooling
Some works add complex modules with attention mechanisms before the conventional pooling strategy. NV-Embed [193] introduces an additional attention-based network before mean pooling to post-interact with the hidden states $\mathcal{H}$. The attention-based network consists of a cross-attention block and an MLP, where the key and value matrix used in the cross-attention are learnable and low-rank, and $\mathcal{H}$ is considered as the query. Mistral-SPLADE [182] uses the prompt same to Echo [185] to obtain text embeddings and follow the SPLADE recipe [198] to fine-tune the Mistral-based embedder.
Multi-Layer Pooling
@@
Some works use post-interactive modules to mix the information from the different layers’ hidden states in the LLM. For example, [194] proposes a trainable multi-layer pooling method that aggregates text embeddings from multiple layers using last/mean pooling. These embeddings form a matrix $\mathcal{H}^{L}$, which is processed by a cross-attention Transformer. Cross-attention is computed by summing $\mathcal{H}^{L}$ with a learnable weight matrix, then deriving key and value representations. A fixed learnable query is used across inputs, and the feed-forward block output serves as the final embedding. [181] collect the hidden states from the layer with the optimal alignment and uniformity metric [199], which denotes $\mathcal{H}^{a}$ and $\mathcal{H}^{u}$ separately. Then, a multi-layer network called LMORT is introduced to fuse the information in both $\mathcal{H}^{a}$ and $\mathcal{H}^{u}$, where each layer consists of a bi-directional self-attention block, a bi-directional cross-attention block, a feed-forward block, and inter-block residual mechanism. The authors show that the performance of retrieval tasks can be improved by only fine-tuning the LMORT’s parameters. MoEE [187] use the routing weights from Mixture-of-Experts (MOE) LLMs as the embeddings, showing a promising complementarity of hidden states from last layers. ReBA [188] aggregates the weights of all attention heads in different layers to recompute the final token embedding and mitigates the shortcomings of causal attention by utilizing a similar “repetition” idea for Echo [188].
4.2.2 Attention Mechanism